how do i determine whether (c+5) is a factor of (c^4+7c^3+6c^2-18c+10)?
Answer by Alan3354(8251)

how do i determine whether (c+5) is a factor of how do i determine whether (c+5) is a factor of (c^4+7c^3+6c^2-18c+10)?
-----------------
The easiest way is to substitute -5 for c. If it's a factor, the polynomial will be zero. (The -5 comes from setting (c+5) = 0.)
(c^4+7c^3+6c^2-18c+10) =? 0
625 -7*125 + 6*25 + 90 + 10 =? 0
625 - 875 + 150 + 90 + 10 =? 0
875 - 875 = 0, so it is a factor.